### v1

@@ Appendix B CTQWs on Network Topologies ranging from Homogeneous to Heterogeneous structures
Figure 10: Quantum-classical distance and fidelity with the initial state for Network size N = 10. Results are shown for different network topologies: cycle, star (with both central and peripheral node initialization), complete, Erdős–Rényi, small-world and scale-free networks (with hub node initialization) under both the noiseless evolution (a),(e) and three decoherence models: Haken–Strobl (b),(f), intrinsic decoherence (c),(g), and the QSW (d),(h).
@@
Figure 11: Time evolution of von Neumann entropy for networks with size N = 10. Von Neumann entropy for all network topologies when the CTQW evolves under (a) Haken–Strobl noise, (b) intrinsic decoherence, and (c) QSW.
Figure 12: Quantum-classical distance and fidelity with the initial state for different network topologies with network size N = 50. The quantum walker is evolved under the conditions: no noise (a),(e), Haken–Strobl noise (b),(f), intrinsic decoherence (c),(g), and the QSW (d),(h). In the quantum-classical distance plots, the gray dashed line for the noiseless case (a) are corresponds to the analytical results.
Figure 13: Von Neumann entropy for Network size N = 50. The quantum walker is evolved under three conditions: Haken–Strobl noise (a), intrinsic decoherence (b), and the QSW (c).
@@
In scale-free networks, nodes often exhibit significant variability in centrality measures such as degree, closeness, and betweenness. We find that when the nodes with the highest degree and highest closeness centrality are distinct, the dynamics of the CTQW exhibit a striking anti-synchronization pattern. Specifically, when the walker is initialized on either the highest-degree node (Fig.14(a)) or the highest-closeness node (Fig.14(b)), the probability distributions over time show out-of-phase oscillations between these two nodes, suggesting a dynamical competition in occupation probabilities.

### v2

@@ Appendix B Extended Stability Metrics for Simple Topologies
Figure 9: Quantum-classical distance and fidelity with the initial state for complete, cycle and star networks with size N = 10. Quantum-classical distance and fidelity with the initial state for noisless (a),(e) and three decoherence models: Haken-Strobl (b),(f), intrinsic Decoherence (c),(g) , and QSW (d),(h). The grey dashed line in (a) is theoretical value of quantum classical distance under no noise
@@
Figure 10: Time evolution of von Neumann entropy for Complete, Cycle and Star networks of size N = 10. Von Neumann entropy under Haken-Strobl (a), Intrinsic Decoherence (b), and QSW (c).
Appendix C Node Centrality Effects in Scale-Free Networks
@@
In scale-free networks, nodes often exhibit significant variability in centrality measures such as degree and closeness. We find that when the nodes with the highest degree and highest closeness centrality are distinct, the dynamics of the CTQW exhibit a striking anti-synchronization pattern. Specifically, when the walker is initialized on either the highest-degree node (Fig.11(a)) or the highest-closeness node (Fig.11(b)), the probability distributions over time show out-of-phase oscillations between these two nodes, suggesting a dynamical competition in occupation probabilities.